# Spatial-Temporal Graph Representation Learning for Tactical Networks Future State Prediction

### v1

@@ I Introduction
To predict future connectivity in the realistic Anglova scenario we introduce a novel Spatial-Temporal Graph Encoder-Decoder (STGED) framework that performs sequence classification on historical network states. Results show that STGED outperform state of the art benchmarks and predict future network connectivity with 99.2% accuracy.
In summary, the main contributions of this paper include:
@@
II Related Works
This section surveys relevant research in TCNs, MANETs, and GNNs, with a particular focus on link prediction and spatial-temporal modeling.
@@ II-C GNNs in Networking
II-D Spatial-Temporal GNNs
Special-Temporal models are widely used to model sequential applications in the natural language processing domain, such as itinerary planning [24], social media mining [25, 26, 27], geo-tagging [28, 29], and other applications [30, 31, 32]. In the domain of communication networks, Spatial-Temporal GNNs (STGNNs) are particularly relevant for capturing data correlation in both time and space. For example, Chen et al. [33] proposed an STGNN combining GCN and long short-term memory (LSTM) unit for cellular network traffic prediction. Wang et al. [34] outperformed traditional GNN and GRU models in predicting cellular traffic. In the context of road traffic forecasting, Lan et al. [35] introduced a dynamic spatial-temporal GNN that significantly outperformed state-of-the-art methods. Li et al. [36] proposed a fusion operation on multiple spatial and temporal graphs for synchronized learning of hidden spatial-temporal correlations. Li et al. [37] proposed a graph encoder-decoder approach to conduct fault diagnosis in industrial processes.
@@ III Methodology
TABLE I: Notations Table
Symbol Description
$TCN$ Tactical Communication Network
$N$ Number of nodes in the TCN
$M$ Number of connected edges in the TCN
$STGED$ The Spatial-Temporal Graph Encoder-Decoder model
$G^{(t)}$ Temporal graph representing the TCN state at time $t$
$V^{(t)}$ Set of node IDs representing the communicating devices at time $t$
$E^{(t)}$ Set of edges indices representing the connectivity at time $t$
$X_{V}^{(t)}$ Node features at time $t$
$X_{E}^{(t)}$ Edge features at time $t$
$C_{X_{V}},C_{X_{W}}$ The number of node features and the number of edge features
$w$ The time window of past network states observation
$t$ The time step of the TCN
$T$ Number of time steps (i.e. states) of the TCN
$SE$ Spatial Encoder unit
$TE$ Temporal Encoder unit
$DE$ Decoder unit
$Z_{(t-w):t}$ Embedding of the TCN from state $t-w$ to state $t$
$H$ Size of the node embedding
@@ III-D Multi-Layer Perceptron Decoder
Figure 2: Node positions over time (in minutes) for the CNTM and CNCM datasets. The top row of figures shows the objective-oriented mobility of the CNTM dataset, the tank company deploys from the upper left to bottom right. The bottom row of figures illustrates the random mobility of the CNCM dataset. Nodes start off in 3 groups and randomly move toward and away from each other.
Figure 3: CNTM t=60m magnified. The tank company travels on a winding road through a hilly area, where the terrain impacts connectivity.
@@ IV-A Datasets
For our TCN experiments we generated datasets using EXata [39], a high fidelity network emulation toolset. EXata uses a software virtual network to represent networks across various protocol levels and radio types, and can even interoperate with real devices to provide hardware-in-the-loop capabilities. We used EXata to model a subset of the Anglova scenario. Table II summarizes the statistics of both datasets.
@@
TABLE II: Summary of statistics of CNTM and CNCM datasets. Note that the Avg. Edges are more than a fully-connected graph with 24 nodes (i.e. 576) because we record the state in one-second interval. Within each time step, there could exist none, one or many edges between any two nodes (i.e. communications between two nodes happen multiple times within one second.)
@@ IV-A3 Visualisation
The differences between the two datasets can be visualized in Figure 2, where we see the random nature of the CNCM dataset and the objective-oriented mobility of the CNTM dataset. Note that the arrows between nodes represent messages being sent successfully. Nodes that are not connected cannot successfully send messages, but not all nodes that are connected will be sending messages at any given snapshot. Also note that the scale is different, as the CNTM travels a greater distance, but generally is grouped together. If the same scale was used for the CNCM dataset, there would be much fewer connections, which would skew the data. To help visualize the CNTM dataset, Figure 3 provides a magnified snapshot of t=60m. In this figure the nodes are on a path that is winding about a large hill, which blocks communication between many nodes on opposite sides. This is one example how the terrain may impact the connectivity more than the distance.
@@ IV-A4 Data Processing
TABLE III: Model Categorization
@@ 1 Step 2 Steps 5 Steps
TABLE IV: Experimental Results for STGED and baselines for the next-state 1-hop connectivity prediction. Using input time steps of one, two and five, STGED consistently outperforms all baseline models across all steps and metrics.
IV-B Baselines
  • GRU [42]: Gated Recurrent Unit, which is a recurrent neural network.
Table III summarises the categories of the models in architectural and encoding types.
IV-C Experiment Setup
@@ IV-D Results
Table IV shows the experimental results of the proposed STGED model and the six baseline methods. Below summarizes our observations and interpretations:
@@
TABLE V: Ablation Experimental Results
@@ IV-E Ablation Experiments
We conduct ablation experiment on proposed STGED framework to variate each component. For spatial encoder, we choose from: (1) GCN, (2) GAT, (3) GATv2 and (4) GTC. For temporal encoder, we choose between LSTM and GRU. All encoder variations will then pass through an MLP as the decoder. Furthermore, we conduct further experimentation to remove the spatial encoder (i.e. with temporal encoder only) or remove the temporal encoder (i.e. with spatial encoder only) to assess their effectiveness. Finally, we use the CNCM dataset to conduct the ablation experiments. As the movements of the nodes are casual without a goal (i.e. highly random), learning an effective representation becomes a challenging task. All models use five time steps as input to predict the 1-hop connectivity of the following state.
Table V shows the results of the ablation experiment. We observe that GTC-LSTM shows the highest accuracy and F1-score. This aligns with our expectation because GTC uses dot-product attention among connected nodes, which has a global vision of latent spatial features; as for the temporal encoder, LSTM consists of more gates (i.e. input, output and forget gates) than the GRU (i.e. reset and update gates). While GRU is time-efficient in training, LSTM is more powerful and adaptable as a temporal encoder.
@@ V Conclusion
In this paper, we present a novel STGED framework for TCN representation learning. We create two datasets to demonstrate the ability of the STGED in learning latent features, such as the spatial-temporal tactical movements. We design experiment tasks to pass through varying network states (i.e. time steps) to the models and predict the next-state one-hop connectivity. Experimental results show that the STGED consistently outperforms all baseline models across all input steps and metrics, revealing advantages in STGED for representing TCNs.

### v2

@@ I Introduction
This work aims to predict future connectivity in the realistic Anglova scenario. In summary, the main contributions include:
@@
II Related Work
@@ II-C GNNs in Networking
II-D Spatial-Temporal Graph Neural Networks (STGNN)
Special-temporal models are widely used to model sequential applications in the natural language processing domain, such as itinerary planning [24], social media mining [25, 26, 27], geo-tagging [28, 29], and other applications [30, 31, 32]. In networking, STGNNs are relevant for capturing data correlation in both time and space. For example, Chen et al. [33] proposed an STGNN combining GCN and long short-term memory (LSTM) unit for cellular network traffic prediction. Wang et al. [34] outperformed traditional GNN and GRU models in predicting cellular traffic. In the context of road traffic forecasting, Lan et al. [35] introduced a dynamic STGNN that significantly outperformed state-of-the-art methods. Li et al. [36] proposed a fusion operation on multiple spatial and temporal graphs for synchronized learning of hidden spatial-temporal correlations. Li et al. [37] proposed a graph encoder-decoder approach to conduct fault diagnosis in industrial processes.
@@ III-D Multi-Layer Perceptron Decoder
Figure 2: Node positions over time (in minutes) for CNTM and CNCM datasets. The top row shows the objective-oriented mobility of the CNTM dataset, the tank company deploys from the upper left to bottom right. The bottom row shows the random mobility of the CNCM dataset. Nodes start off in 3 groups and randomly move toward and away from each other.
Figure 3: CNTM t=60m magnified
@@ IV-A Datasets
For our TCN experiments we generated datasets using EXata [39], a high fidelity network emulation toolset. EXata uses a software virtual network to represent networks across various protocol levels and radio types, and can even interoperate with real devices to provide hardware-in-the-loop capabilities. We used EXata to model a subset of the Anglova scenario. Table I summarizes the statistics of both datasets.
@@
TABLE I: Statistics of CNTM and CNCM datasets. Note that the Avg. Edges are more than a fully-connected graph with 24 nodes in one-second interval. Within each time step, there could exist none, one or many edges between any two nodes.
@@ IV-A3 Visualisation
As shown in Figure 2, the random nature of the CNCM dataset and the objective-oriented mobility of the CNTM dataset. Note that the arrows between nodes represent messages being sent successfully. Nodes that are not connected cannot successfully send messages, but not all nodes that are connected will be sending messages at any given snapshot. Also note that the scale is different, as the CNTM travels a greater distance, but generally is grouped together. If the same scale was used for the CNCM dataset, there would be much fewer connections, which would skew the data. To help visualize the CNTM dataset, Figure 3 provides a magnified snapshot of t=60m. In this figure the nodes are on a path that is winding about a large hill, which blocks communication between many nodes on opposite sides. This is one example how the terrain may impact the connectivity more than the distance.
@@ IV-A4 Data Processing
TABLE II: Model Categorization
IV-B Baselines
  • GRU [42]: Gated Recurrent Unit, which is a recurrent neural network.
Table II summarises the categories of the models in architectural and encoding types.
IV-C Experiment Setup
@@ 1 Step 2 Steps 5 Steps
TABLE III: Experimental Results for STGED and baselines for the next-state 1-hop connectivity prediction. Using input time steps of one, two and five, STGED consistently outperforms all baseline models across all steps and metrics.
@@ IV-D Results
Table III shows the experimental results of the proposed STGED model and the six baseline methods. Below summarizes our observations and interpretations:
@@
TABLE IV: Ablation Experimental Results
@@ IV-E Ablation Experiments
We conduct ablation experiment on proposed STGED framework to variate each component. For spatial encoder, we choose from: (1) GCN, (2) GAT, (3) GATv2 and (4) GTC. For temporal encoder, we choose between LSTM and GRU. Furthermore, we conduct further experimentations to assess effectiveness of temporal-encoder-only and spatial-encoder-only models. Finally, we use the CNCM dataset, i.e., random movements of the nodes without a goal, to benchmark the effectiveness in representation learning. All models use five time steps as input to predict the 1-hop connectivity of the following state.
Table IV shows the results of the ablation experiment. We observe that GTC-LSTM shows the highest accuracy and F1-score. This aligns with our expectation because GTC uses dot-product attention among connected nodes, which has a global vision of latent spatial features; as for the temporal encoder, LSTM consists of more gates (i.e. input, output and forget gates) than the GRU (i.e. reset and update gates). While GRU is time-efficient in training, LSTM is more powerful and adaptable as a temporal encoder.
@@ V Conclusion
This paper presents a novel STGED framework for TCN representation learning. Two datasets demonstrate the ability of the STGED in learning latent features, such as the spatial-temporal tactical movements. We design experiment tasks to pass through varying network states (i.e. time steps) to the models and predict the next-state one-hop connectivity. Experimental results show that the STGED consistently outperforms all baseline models across all input steps and metrics, revealing advantages in STGED for representing TCNs.
Acknowledgement
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Future Communications and Research & Development Programme.